What is X2+54X+10


1
Expert's answer
2021-02-09T07:07:44-0500

using quadratic equation

x= "\frac{-b\frac{+}{} \sqrt(4ac)}{2a}"

="\frac{-54\frac{+}{} \sqrt(54*54 -4*10)}{2}"

= "\frac{-54 \frac{+}{}{53.628}}{2},"

= -0.158 or -53.8